OpenStudy (anonymous):

given this equation: y1 x y'2 - y2 x y'1 = ? y1= e^(-2t) y2=e^(-2t) What is the answer? I keep getting 0...

OpenStudy (anonymous):

e^(-2t) x (-2e^(-2t)) - e^(-2t) x (-2e^(-2t)) -2e^(4t)+2e^(4t)=0 am i right?

OpenStudy (anonymous):

yeah you are right because look y1=y2 so the derivatives will be the same as well right so yes if you subtract them you will get zero :)

OpenStudy (anonymous):

but it will be e^(-4t) not e^4t alright

OpenStudy (anonymous):

ok thanks man!

OpenStudy (anonymous):

even when you multiply e^(-2t) x e^(-2t) the multiplication by negatives doesn't make them positive?

OpenStudy (anonymous):

if you are multiplying exponent then you must add the top so e^(-2t - 2t) = e^(-4t) :)

OpenStudy (anonymous):

ok perfect thanks again